The car needs to jump through horizontal distance of at least 20 Cars with some suitable height. The Car driver wishes to take the complete one role manoeuvre i.e. rotation along the motion of axis, into the air so that manoeuvre be completed successfully (for rigid body-based analysis condition).
1
Expert's answer
2020-12-21T13:40:19-0500

Answer

Using energy conservation here v'=0

"\frac{mv'^2}{2}+mgy=\frac{mv^2}{2}+mg\Delta y"

So velocity is become

"v=\sqrt{v'^2+2g(y-\Delta y) }"

Using energy conservation here v''=0

"\frac{mv^2}{2}+mg\Delta y_2=\frac{mv''^2}{2}+mgH_m"

Using above result

"H_m=\Delta y_2+(\frac{v''^2}{2g}+(y-\Delta y)) \cos\theta"